OpenStudy (anonymous):

Find f(x-3) when f(x)=x-3

OpenStudy (anonymous):

f(x-3) = (x-3) -3 = x-6

hero (hero):

You always start with f(x) f(x) = x - 3 The x - 3 part is then inserted in place of x: f(x - 3) = x - 3 - 3 then you simplify to get: f(x - 3) = x - 6

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

what about f(x+3)? would that just be x?

hero (hero):

By default, f(x) will always be your original function, so if you see f(x+3), you know that x was replaced with x + 3

OpenStudy (anonymous):

ok